Circular No. 7089

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999E IN ANONYMOUS GALAXY I. Perez and J. Maza, Department of Astronomy, University of Chile,

reports the discovery of a supernova (B ∼ 18) by Roberto Antezana on a T-Max 400 film taken by L. E. González using the Maksutov telescope at Cerro El Roble on Jan. 15.276 UT. Direct CCD images by Perez on Jan. 16.374 using the 1.0-m Swope telescope at Las Campanas Observatory yield V ∼ 16. SN 1999E is located at α = 13^{h}17^{m}16^{s}.37, δ = −18^{o}33′13′′.4 (equinox 2000.0), which is 0′′.9 west and 10′′.0 south of the nucleus of the host spiral galaxy. The new star is not visible on a film obtained on 1998 July 29.046 (limiting mag B > 20).

SUPERNOVA 1999D IN NGC 3690 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1999D, obtained by M. Calkins on Jan. 18.5 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova at an early epoch. The spectrum features a very blue continuum, with strong P-Cyg Balmer lines. Narrow Hα emission indicates a recession velocity for NGC 3690 of 2900 km/s. The photospheric expansion velocity of the supernova is 11 000 km/s, measured to the Hα trough. Concurrent images taken by D. Koranyi at the Whipple Observatory 1.2-m telescope yield a position for the supernova slightly different from that reported by Qiu et al. (IAUC 7088). The supernova appears at α = 11^{h}28^{m}28^{s}.38, δ = +58^{o}33′39′′.0 (equinox 2000.0, uncertainty ± 0′′.2), which is 18′′.9 west and 4′′.8 south of the nucleus of NGC 3690. The spectrum and a finder chart can be obtained at http://cfa-

www.harvard.edu/cfa/oir/Research/supernova/RecentSN.html.

SUPERNOVA 1999B IN UGC 7189 A. V. Filippenko, D. C. Leonard, and A. G. Riess, University of Cali-

fornia at Berkeley, report that a Keck-II spectrum (range 540–910 nm) of SN 1999B obtained on Jan. 18 UT confirms that this is a type-II supernova (cf. IAUC 7088). The spectrum exhibits strong, broad Hα emission, and a few weak absorption lines.

1999 January 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green